Hi.
What is the least common denominator of the following set:
1/(a+1)^2, 1/(a-1)^2, and 2/a^2-1

Thanks for your help!

Found 2 solutions by stanbon, gems:
Answer by stanbon(75887)

What is the least common denominator of the following set:
1/(a+1)^2, 1/(a-1)^2, and 2/a^2-1
--------------------
It is (a+1)^2*(a-1)^2 or (a^2-1)^2
==========================
Cheers,
Stan H.

Answer by gems(12)

the factors of (a+1)^2 = ( a+1)(a+1)
the factors of (a-1)^2= (a-1)(a-1)
the factors of a^2-1 = (a+1)(a-1)
therefore the LCD = ( a+1)(a+1)(a-1)(a-1)
or (a+1)^2(a-1)^2